ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 140, 2004

Bar-Natan, D. and R. Lawrence A rational surgery formula for the LMO invariant

29

Bujalance, J. A., A. F. Costa and A. M. Porto On the topological types of symmetries of elliptic-hyperelliptic Riemann surfaces

145

Castillo, J. M. F. and Y. Moreno On the Lindenstrauss– Rosenthal theorem

253

Costa, A. F. See J. A. Bujalance

Duda, J. On the size of the set of points where the metric projection exists

271

Gao, W. D. and D. J. Wang Additive Latin transversals and group rings

375

Gorodnik, A. On an Oppenheim-type conjecture for systems of quadratic forms

125

Inoue, K. Isometric extensions and multiple recurrence of infinite measure preserving systems

245

Kamińska, A. and L. Maligranda On Lorentz spaces Γ_{p,w}  285

Lapid, E. and S. Rallis A spectral identity between symmetric spaces

221

Lawrence, R. See D. Bar-Natan

Maligranda, L. See A. Kamińska

Matsuzaki, K. Inclusion relations between the Bers embeddings of Teichmüller spaces

113

Mendelson, S. and R. Vershynin Remarks on the geometry of coordinate projections in ℝ^{n}

203

Moreno, Y. See J. M. F. Castillo

Muić, G. Composition series of generalized principal series; the case of strongly positive discrete series

157

Porto, A. M. See J. A. Bujalance

Rallis, S. See E. Lapid

Randrianantoanina, N. Square function inequalities for noncommutative martingales 333 Shelah, S. Characterizing an ℵ_{ϵ}-saturated model of superstable

NDOP theories by its 𝕃_{∞,ℵ_{ϵ}}-theory

61

Skorobogatov, A. and A. Yafaev Descent on certain Shimura curves

319

Temkin, M. On local properties of non-Archimedean analytic spaces II

1

Tsaban, B. The minimal cardinality where the Reznichenko property fails

367

Vershynin, R. See S. Mendelson

Wang, D. J. See W. D. Gao

Yafaev, A. See A. Skorobogatov